Question 1

On the first day of its fiscal year, Chin Company issued $26,500,000 of five-year, 9% bonds to finance its operations of producing and selling home improvement products. Interest is payable semiannually. The bonds were issued at a market (effective) interest rate of 11%, resulting in Chin Company receiving cash of $24,502,519.

a. Journalize the entries to record the following:

  1. Issuance of the bonds.
  2. First semiannual interest payment. The bond discount amortization, using the straight-line method, is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)
  3. Second semiannual interest payment. The bond discount amortization, using the straight-line method, is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)

b. Determine the amount of the bond interest expense for the first year.
$

c. Why was the company able to issue the bonds for only $24,502,519 rather than for the face amount of $26,500,000?
The market rate of interest is the contract rate of interest.

Question 2

Daan Corporation wholesales repair products to equipment manufacturers. On April 1, 2016, Daan Corporation issued $4,900,000 of 5-year, 8% bonds at a market (effective) interest rate of 5%, receiving cash of $5,543,276. Interest is payable semiannually on April 1 and October 1.

a. Journalize the entry to record the issuance of bonds on April 1, 2016. For a compound transaction, if an amount box does not require an entry, leave it blank.

b. Journalize the entry to record the first interest payment on October 1, 2016, and amortization of bond premium for six months, using the straight-line method. The bond premium amortization is combined with the semiannual interest payment. (Round to the nearest dollar.) For a compound transaction, if an amount box does not require an entry, leave it blank.

c. Why was the company able to issue the bonds for $5,543,276 rather than for the face amount of $4,900,000?

The market rate of interest is the contract rate of interest.
Answer #1
Question 1
Face Value of Bond $                                                       2,65,00,000
Issue Price of Bond $                                                       2,45,02,519
Discount on issue of Bond $                                                          19,97,481
Amortization of Discount on Bond over 10 period =$1,997,481 / 10 =$199,748
Interest paid on Bond =$26,500,000*9%96/12 =$1,192,500
Interest expenses on Bond =$1,192,500 + $199,748 =$1,392,248
Date Accounts and explanation Debit(in $) Credit(in $)
Issue date Cash $                                                       2,45,02,519
Discount on Bonds Payable $                                                          19,97,481
Bond Payable $                            2,65,00,000
First semiannual interest payment Interest expenses $                                                          13,92,248
Cash $                               11,92,500
Discount on Bonds Payable $                                 1,99,748
Second semiannual interest payment Interest expenses $                                                          13,92,248
Cash $                               11,92,500
Discount on Bonds Payable $                                 1,99,748
Bond Interest expense for Year 1 =$1,392,248*2 =$2,784,496
The Company received only $24,502,519 because the market interest rate is greater than the stated interest rate and that is why bond is issued at discount
Question 2
Issue Price of Bond $                                                          55,43,276
Face Value of Bond $                                                          49,00,000
Premium on issue of Bond $                                                            6,43,276
Amortization of Discount on Bond over 10 period =$643,276/10 =$64,328
Interest paid on Bond =$4,900,000*8%*6/12 =$196,000
Interest expenses on Bond =$196,000 - $64,328 =$131,672
Date Accounts and explanation Debit(in $) Credit(in $)
Issue date Cash $                                                          55,43,276
Premium on Bond payable $                                 6,43,276
Bond Payable $                               49,00,000
First semiannual interest payment Interest expenses $                                                            1,31,672
Premium on Bond payable $                                                               64,328
Cash $                                 1,96,000
Second semiannual interest payment Interest expenses $                                                            1,31,672
Premium on Bond payable $                                                               64,328
Cash $                                 1,96,000
Bond Interest expense for Year 1 =$131,672*2 =$263,344
The Company received $5,543,276 because the market interest rate is lower than the stated interest rate and that is why bond is issued at premium